# Propagation of Chaos for Mean-Field Langevin Dynamics and its Application to Model Ensemble

### v1

@@ 5.2 LoRA for finetuning language models
The first equality of the assertion was proved by Nitanda (2024). We here prove the inequality by utilizing the argument of conditional and marginal distribution of $\mu^{(N)}$ Chen et al. (2022).
For $\mathbf{X}\sim\mu^{(N)}$, we denote by $\mu^{(N)}_{i|-i}(\cdot|\mathbf{x}^{-i})$ and $\mu^{(N)}_{-i}$ the conditional distribution of $X^{i}$ conditioned by $\mathbf{X}^{-i}=\mathbf{x}^{-i}$ and the marginal distribution of $\mathbf{X}^{-i}$, respectively. It holds that
@@
$=\sum_{i=1}^{N}\mathbb{E}_{\mathbf{x}\sim\mu^{(N)}}\left[\left\|\nabla_{x^{i}}\log\frac{\mathrm{d}\mu^{(N)}}{\mathrm{d}\mathbf{x}}(\mathbf{X})+\frac{N}{\lambda}\nabla_{x^{i}}F(\mu_{\mathbf{x}})\right\|_{2}^{2}\right]$
$=\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}^{-i}\sim\mu^{(N)}_{-i}}\left[\mathbb{E}_{X^{i}\sim\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})}\left[\left\|\nabla_{x^{i}}\log\frac{\mathrm{d}\mu^{(N)}}{\mathrm{d}\mathbf{x}}(\mathbf{X})+\frac{N}{\lambda}\nabla_{x^{i}}F(\mu_{\mathbf{x}})\right\|_{2}^{2}\right]\right].$ (10)
@@
$\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}^{-i}\sim\mu^{(N)}_{-i}}\left[\mathbb{E}_{X^{i}\sim\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})}\left[\left\|\nabla\log p_{i|-i}(X^{i}|\mathbf{X}^{-i})+\frac{N}{\lambda}\nabla_{x^{i}}F(\mu_{\mathbf{x}})\right\|_{2}^{2}\right]\right]$
@@
Let $\nu$ be the probability distribution on $\mathbb{R}^{d}$ with the density $\frac{\mathrm{d}\nu}{\mathrm{d}x}(x)\propto\exp(-r(x)/\lambda)$. Here, notice that the conditional Gibbs distribution $\nu_{i|-i}(\cdot|\mathbf{x}^{-i})$ is the minimizer of the following objective: for $\mathbf{x}^{-i}$
$\nu_{i|-i}(\cdot|\mathbf{x}^{-i})$ $=\mathop{\mathrm{argmin}}_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}\left\{\int NF_{0}(\rho_{x\cup\mathbf{x}^{-i}})\mu(\mathrm{d}x)+\mathbb{E}_{X\sim\mu}[r(X)]+\lambda\mathrm{Ent}(\mu)\right\}$
$=\mathop{\mathrm{argmin}}_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}\left\{\int NF_{0}(\rho_{x\cup\mathbf{x}^{-i}})\mu(\mathrm{d}x)+\lambda\mathrm{KL}(\mu\|\nu)\right\}.$
Because of the optimality of the conditional Gibbs distribution, we have
$\lambda\mathrm{KL}(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})\|\nu_{i|-i}(\cdot|\mathbf{X}^{-i}))$
$=\lambda\mathrm{Ent}(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i}))+\int NF(\rho_{x\cup\mathbf{X}^{-i}})\mu^{(N)}_{i|-i}(\mathrm{d}x|\mathbf{X}^{-i})+\lambda\log\int\exp\left(-\frac{N}{\lambda}F(\rho_{x\cup\mathbf{X}^{-i}})\right)\mathrm{d}x$
$=\lambda\mathrm{KL}\left(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})\|\nu\right)+\int NF_{0}(\rho_{x\cup\mathbf{X}^{-i}})\left(\mu^{(N)}_{i|-i}(\mathrm{d}x|\mathbf{X}^{-i})-\nu_{i|-i}(\mathrm{d}x|\mathbf{X}^{-i})(\mathrm{d}x)\right)-\lambda\mathrm{KL}\left(\nu_{i|-i}(\cdot|\mathbf{X}^{-i})\|\nu\right)$
$\geq\lambda\mathrm{KL}\left(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})\|\nu\right)+\int NF_{0}(\rho_{x\cup\mathbf{X}^{-i}})\left(\mu^{(N)}_{i|-i}(\mathrm{d}x|\mathbf{X}^{-i})-\mu_{*}(\mathrm{d}x)\right)-\lambda\mathrm{KL}\left(\mu_{*}\|\nu\right).$ (12)
The expectation of the second term of the right-hand side can be futher evaluated as
$N\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\int F_{0}(\rho_{x\cup\mathbf{X}^{-i}})\left(\mu^{(N)}_{i|-i}(\mathrm{d}x|\mathbf{X}^{-i})-\mu_{*}(\mathrm{d}x)\right)\right]$
$=N\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[F_{0}(\rho_{\mathbf{X}})-\int F_{0}(\rho_{x\cup\mathbf{X}^{-i}})\mu_{*}(\mathrm{d}x)\right]$
$=-N\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\int\left\{B_{F_{0}}(\rho_{x\cup\mathbf{X}^{-i}},\rho_{\mathbf{X}})+\left\langle\frac{\delta F_{0}(\rho_{\mathbf{X}})}{\delta\mu},\rho_{x\cup\mathbf{X}^{-i}}-\rho_{\mathbf{X}}\right\rangle\right\}\mu_{*}(\mathrm{d}x)\right]$
$=-N\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\int B_{F_{0}}(\rho_{x\cup\mathbf{X}^{-i}},\rho_{\mathbf{X}})\mu_{*}(\mathrm{d}x)\right]$
$~{}~{}~{}~{}+\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\int\left\{\frac{\delta F_{0}(\rho_{\mathbf{X}})}{\delta\mu}(X^{i})-\frac{\delta F_{0}(\rho_{\mathbf{X}})}{\delta\mu}(x)\right\}\mu_{*}(\mathrm{d}x)\right]$
$=-N\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\int B_{F_{0}}(\rho_{x\cup\mathbf{X}^{-i}},\rho_{\mathbf{X}})\mu_{*}(\mathrm{d}x)\right]$
$~{}~{}~{}~{}+N\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\int\frac{\delta F_{0}(\rho_{\mathbf{X}})}{\delta\mu}(x)\left(\rho_{\mathbf{X}}(\mathrm{d}x)-\mu_{*}(\mathrm{d}x)\right)\right]$
$\geq-B+N\left(\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[F_{0}(\rho_{\mathbf{X}})\right]-F_{0}(\mu_{*})\right),$ (13)
where the last inequality is due to the convexity of $F_{0}$ and Assumption 4.
By the information inequality (Lemma 5.1 of Chen et al. (2022)), the first term of Eq. (12) of the right-hand side can be evaluated as
$\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\mathrm{KL}\left(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})\|\nu\right)\right]\geq\mathrm{KL}(\mu^{(N)}\|\nu^{\otimes N}).$ (14)
Combining all of them, we get
$\geq\frac{2\alpha}{\lambda}\left\{-B+N\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[F_{0}(\rho_{\mathbf{X}})\right]+\lambda\mathrm{KL}(\mu^{(N)}\|\nu^{\otimes N})-N(F_{0}(\mu_{*})+\lambda\mathrm{KL}\left(\mu_{*}\|\nu\right))\right\}$
$=\frac{2\alpha}{\lambda}\left\{-B+N\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[F(\rho_{\mathbf{X}})\right]+\lambda\mathrm{Ent}(\mu^{(N)})-N(F(\mu_{*})+\lambda\mathrm{Ent}(\mu_{*}))\right\}$
$=\frac{2\alpha}{\lambda}\left(-B+\mathcal{L}^{(N)}(\mu^{(N)})-N\mathcal{L}(\mu_{*})\right).$
@@
We here prove the convergence of MFLD in the discrete-time by using the one-step interpolation argument Nitanda (2024); Suzuki et al. (2023a).
@@
By the proof of Theorem 2 in Nitanda (2024), we see for $t\in[0,\eta]$,
@@
[…]
@@
It follows that by Han’s inequality (Dembo et al., 1991),
@@
We evaluate the empirical Rademacher complexity ^$\hat{\mathcal{R}}_{N,M}(\mathcal{F})$ by using Dudley’s entropy integral. We define the metric $\|f\|_{N,M,2}=\sqrt{\frac{1}{MN}\sum_{j=1}^{M}\sum_{i=1}^{N}|f(X^{i}_{j})|^{2}}$. We denote by $\mathcal{N}(\mathcal{F},\epsilon,\|\cdot\|_{N,M,2})$ the $\epsilon$-covering number of $\mathcal{F}$ with respect to the $\|\cdot\|_{N,M,2}$-norm.
@@
^$\hat{\mathcal{R}}_{N,M}(\mathcal{F})\leq\inf_{\delta>0}\left\{4\delta+\frac{12}{\sqrt{MN}}\int_{\delta}^{R}\sqrt{\log 2\mathcal{N}(\mathcal{F},\epsilon,\|\cdot\|_{N,M,2})}\mathrm{d}\epsilon\right\}.$
@@
^$\hat{\mathcal{R}}_{N,M}(\mathcal{F})$ $\leq 4R\sqrt{\frac{d}{MN}}+12R\sqrt{\frac{1}{MN}\log 2\mathcal{N}(\mathcal{F},R\sqrt{d(MN)^{-1}},\|\cdot\|_{N,M,\infty})}$
@@
For $\mathbf{x}_{1},\ldots,\mathbf{x}_{M}\in\mathbb{R}^{dN}$, we set $g(\mathbf{x}_{1},\ldots,\mathbf{x}_{M})=\sup_{z\in\mathcal{Z}}\left|\frac{1}{M}\sum_{j=1}^{M}\mathbb{E}_{X\sim\rho_{\mathbf{x}_{j}}}[h(X,z)]-\mathbb{E}_{X\sim\mu_{*}}[h(X,z)]\right|$. By the variational formulation of KL-divergence (e.g., Corollary 4.15 in Boucheron et al. (2013)), we get
@@
0:  $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}=(z_{\text{test}},y_{\text{test}})$, $N_{\infty}$, $N_{\text{list}}$, $M_{\text{max}}$
0:  Dictionary sup_norm_dic maps $N$ to the average sup_norm
@@
     Initialize sup_norm_lst [$\leftarrow$[]← [ ]
@@
           sup_norm $\leftarrow\text{max}\left(|\hat{y}-\hat{y}_{\infty}|\right)$
@@
Figure 2: Averaged test ln(MSE) of singular MFNNs, across different $N$ and $\lambda$ for 5 epochs
Table 2: MSE comparison between merging $M=20$ networks across different $N$ and $\lambda$ after 5 epochs.
@@
Llama2 7B LoRA (best) 00 $80.55$ $82.86$ $83.19$ $81.60$ $71.08$ $84.51$ $71.90$ $90.21$ $80.74$
LoRA (ave.) 00 $64.73$ $76.31$ $77.76$ $68.70$ $57.02$ $69.02$ $69.04$ $70.63$ $69.15$
PoC merge 00 $32.29$ $62.57$ $83.58$ $22.20$ $28.41$ $29.42$ $61.53$ $28.50$ $43.56$
@@
Llama3 8B LoRA (best) 00 $80.45$ $88.47$ $86.82$ $87.60$ $82.25$ $90.87$ $73.85$ $95.78$ $85.76$
LoRA (ave.) 00 $80.51$ $88.87$ $86.85$ $87.00$ $80.78$ $90.98$ $73.71$ $95.84$ $85.57$
PoC merge 00 $81.73$ $88.96$ $87.77$ $88.00$ $81.40$ $91.71$ $74.46$ $96.45$ $86.31$
@@
Two layer mean-field neural networks provide a tractable analytical framework for studying infinitely wide neural networks. As $N\rightarrow\infty$, the optimization dynamics is captured by a partial differential equation (PDE) of the parameter distribution, where convexity can be exploited to show convergence to the global optimal solution (Chizat & Bach, 2018; Mei et al., 2018; Rotskoff et al., 2019). If Gaussian noise is added to the gradient, we get MFLD which achieves global convergence to the optimal solution (Mei et al., 2018; Hu et al., 2019). Under the uniform LSI, Nitanda et al. (2022); Chizat (2022) show that MFLD converges at an exponential rate by using the proximal Gibbs distribution associated with the dynamics. MFLD has attracted significant attention because of its feature learning capabilities (Suzuki et al., 2023b; Mousavi-Hosseini et al., 2024).
As the assumption that $N=\infty$ is not applicable to real-world scenarios, a discrete-time finite particle system would align closer to an implementable MFLD i.e. noisy gradient descent. Nitanda et al. (2022) provides a global convergence rate analysis for the discrete-time update by extending the one-step interpolation argument for Langevin dynamics (Vempala & Wibisono, 2019). Meanwhile, the approximation error of the finite particle setting is studied in propagation of chaos literature (Sznitman, 1991). For finite MFLD setting, Mei et al. (2018) first suggested that approximation error grows exponentially with time before Chen et al. (2022); Suzuki et al. (2023a) proved the uniform-in-time propagation of chaos with error bound: $O\left(\frac{\alpha}{\lambda N}\right)$, suggesting that the difference between the finite $N$-particle system and mean-field limit shrinks as $N\rightarrow\infty$. However, this also means that particle approximation error blows-up exponentially as $\lambda\rightarrow 0$ due to the exponential relationship between $\alpha$ and $\lambda$ (Nitanda et al., 2022; Chizat, 2022; Suzuki et al., 2023a). Suzuki et al. (2023b) proposes an annealing procedure for classification tasks to remove this exponential dependence in LSI, wihch requires that $\lambda$ be gradually reduced over time and will not work for fixed regularization parameters. Nitanda (2024); Chewi et al. (2024) then proved a refined propagation of chaos independent of $\alpha$ at the solution as described in Section 1.1.

### v2

@@ Acknowledgements
The first equality of the assertion was proved by Nitanda, (2024). We here prove the inequality by utilizing the argument of the conditional and marginal distribution of $\mu^{(N)}$ Chen et al., (2022).
By the convexity, we have that for any $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
$\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[F(\rho_{\mathbf{X}})+\left\langle\frac{\delta F(\rho_{\mathbf{X}})}{\delta\mu},\mu-\rho_{\mathbf{X}}\right\rangle\right]+\lambda\mathrm{Ent}(\mu)\leq\mathcal{L}(\mu).$ (11)
To further evaluate the lower-bound, we begin with the following equation.
$\lambda\log Z(\mathbf{x}^{-i})=\min_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}\left\{N\mathbb{E}_{X\sim\mu}[F(\rho_{X\cup\mathbf{x}^{-i}})]+\lambda\mathrm{Ent}(\mu)\right\},$ (12)
where $Z(\mathbf{x}^{-i})=\int\exp\left(-\frac{N}{\lambda}F(\rho_{x\cup\mathbf{x}^{-i}})\right)\mathrm{d}x$ is the normalizing constant of $\nu_{i|-i}(\cdot|\mathbf{x}^{-i})$.
This equality is confirmed by reformulating the objective of (12) as follows:
$N\mathbb{E}_{X\sim\mu}[F(\rho_{x\cup\mathbf{X}^{-i}})]+\lambda\mathrm{Ent}(\mu)=\lambda\mathrm{KL}(\mu\|\nu_{i|-i}(\cdot|\mathbf{X}^{-i}))-\lambda\log Z(\mathbf{X}^{-i}).$
Then, the lower-bound in Eq. (11) can be evaluated as follows:
$\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[F(\rho_{\mathbf{X}})+\left\langle\frac{\delta F(\rho_{\mathbf{X}})}{\delta\mu},\mu-\rho_{\mathbf{X}}\right\rangle\right]+\lambda\mathrm{Ent}(\mu)$
$=\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[F(\rho_{\mathbf{X}})+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{X\sim\mu}\left[\frac{\delta F(\rho_{\mathbf{X}})}{\delta\mu}(X)-\frac{\delta F(\rho_{\mathbf{X}})}{\delta\mu}(X^{i})\right]\right]+\lambda\mathrm{Ent}(\mu)$
$=\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\mathbb{E}_{X\sim\mu}\left[F(\rho_{\mathbf{X}})+\left\langle\frac{\delta F(\rho_{\mathbf{X}})}{\delta\mu},\rho_{X\cup\mathbf{X}^{-i}}-\rho_{\mathbf{X}}\right\rangle\right]-(N-1)\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}[\mathcal{F}(\rho_{\mathbf{X}})]+\lambda\mathrm{Ent}(\mu)$
$=\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\mathbb{E}_{X\sim\mu}\left[F(\rho_{X\cup\mathbf{X}^{-i}})-B_{F}(\rho_{X\cup\mathbf{X}^{-i}},\rho_{\mathbf{X}})\right]-(N-1)\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}[\mathcal{F}(\rho_{\mathbf{X}})]+\lambda\mathrm{Ent}(\mu)$
$\geq-\frac{\lambda}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}[\log Z(\mathbf{X}^{-i})]-\frac{B}{N}-(N-1)\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}[\mathcal{F}(\rho_{\mathbf{X}})],$ (13)
where we used Assumption 4 for the last inequality.
We consider the following decomposition:
$\frac{\lambda}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}$ $\left[\mathrm{KL}(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})\|\nu_{i|-i}(\cdot|\mathbf{X}^{-i}))\right]$
$=\frac{\lambda}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\mathrm{Ent}(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i}))\right]$
$+\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\int F(\rho_{x\cup\mathbf{X}^{-i}})\mu^{(N)}_{i|-i}(\mathrm{d}x|\mathbf{X}^{-i})\right]$
$+\frac{\lambda}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\log Z(\mathbf{X}^{-i})\right].$ (14)
The first term can be bounded by the well-known property of entropy.
$\frac{\lambda}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\mathrm{Ent}(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i}))\right]\geq\frac{\lambda}{N}\mathrm{Ent}(\mu^{(N)}).$ (15)
The second term can be simply written as follows.
$\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\int F(\rho_{x\cup\mathbf{X}^{-i}})\mu^{(N)}_{i|-i}(\mathrm{d}x|\mathbf{X}^{-i})\right]=N\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}[F(\rho_{\mathbf{X}})].$ (16)
Combining (A. 1) (A. 1), (16), and (15), we get
$\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[F(\rho_{\mathbf{X}})+\left\langle\frac{\delta F(\rho_{\mathbf{X}})}{\delta\mu},\mu-\rho_{\mathbf{X}}\right\rangle\right]+\lambda\mathrm{Ent}(\mu)$
$\geq-\frac{\lambda}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\mathrm{KL}(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})\|\nu_{i|-i}(\cdot|\mathbf{X}^{-i}))\right]+\frac{\lambda}{N}\mathrm{Ent}(\mu^{(N)})-\frac{B}{N}+\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}[\mathcal{F}(\rho_{\mathbf{X}})]$
$\geq-\frac{B}{N}+\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)})-\frac{\lambda}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\mathrm{KL}(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})\|\nu_{i|-i}(\cdot|\mathbf{X}^{-i}))\right].$ (17)
This concludes
$\frac{\lambda}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\mathrm{KL}(\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})\|\nu_{i|-i}(\cdot|\mathbf{X}^{-i}))\right]\geq-\frac{B}{N}+\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)})-\mathcal{L}(\mu).$
We finish the proof by setting $\mu=\mu_{*}$. ∎
Proof of Lemma 2.
We denote by $\mu^{(N)}_{-i}$ the marginal distribution over $\mathbf{X}^{-i}$. It holds that
@@
$=\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}\sim\mu^{(N)}}\left[\left\|\nabla_{x^{i}}\log\frac{\mathrm{d}\mu^{(N)}}{\mathrm{d}\mathbf{x}}(\mathbf{X})+\frac{N}{\lambda}\nabla_{x^{i}}F(\rho_{\mathbf{X}})\right\|_{2}^{2}\right]$
$=\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}^{-i}\sim\mu^{(N)}_{-i}}\left[\mathbb{E}_{X^{i}\sim\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})}\left[\left\|\nabla_{x^{i}}\log\frac{\mathrm{d}\mu^{(N)}}{\mathrm{d}\mathbf{x}}(\mathbf{X})+\frac{N}{\lambda}\nabla_{x^{i}}F(\rho_{\mathbf{X}})\right\|_{2}^{2}\right]\right].$ (18)
@@
$\sum_{i=1}^{N}\mathbb{E}_{\mathbf{X}^{-i}\sim\mu^{(N)}_{-i}}\left[\mathbb{E}_{X^{i}\sim\mu^{(N)}_{i|-i}(\cdot|\mathbf{X}^{-i})}\left[\left\|\nabla\log p_{i|-i}(X^{i}|\mathbf{X}^{-i})+\frac{N}{\lambda}\nabla_{x^{i}}F(\rho_{\mathbf{X}})\right\|_{2}^{2}\right]\right]$
@@
Combining this inequality and Lemma 1, we get
$\mathbb{E}_{\mathbf{x}\sim\mu^{(N)}}\left[\left\|\nabla\log\frac{\mathrm{d}\mu^{(N)}}{\mathrm{d}\mu^{(N)}_{*}}(\mathbf{X})\right\|_{2}^{2}\right]\geq\frac{2\alpha}{\lambda}\left(-B+\mathcal{L}^{(N)}(\mu^{(N)})-N\mathcal{L}(\mu_{*})\right).$
@@
We here prove the convergence of MFLD in the discrete-time by using the one-step interpolation argument Nitanda, (2024); Suzuki et al., 2023a .
@@
By the proof of Theorem 2 in Nitanda, (2024), we see for $t\in[0,\eta]$,
@@
$\Longleftrightarrow$~~~~ $\frac{\mathrm{d}}{\mathrm{d}t}\left(\mathcal{L}^{(N)}(\nu_{t})-N\mathcal{L}(\mu_{*})-B-\frac{N\delta_{\eta}}{\alpha\lambda}\right)\leq-\alpha\lambda\left(\mathcal{L}^{(N)}(\nu_{t})-N\mathcal{L}(\mu_{*})-B-\frac{N\delta_{\eta}}{\alpha\lambda}\right).$
@@
It follows that by Han’s inequality (Dembo et al.,, 1991),
@@
We evaluate the empirical Rademacher complexity $\hat{\mathcal{R}}_{N,M}(\mathcal{F})$ by using Dudley’s entropy integral. We define the metric $\|f\|_{N,M,2}=\sqrt{\frac{1}{MN}\sum_{j=1}^{M}\sum_{i=1}^{N}|f(X^{i}_{j})|^{2}}$. We denote by $\mathcal{N}(\mathcal{F},\epsilon,\|\cdot\|_{N,M,2})$ the $\epsilon$-covering number of $\mathcal{F}$ with respect to the $\|\cdot\|_{N,M,2}$-norm.
@@
$\hat{\mathcal{R}}_{N,M}(\mathcal{F})\leq\inf_{\delta>0}\left\{4\delta+\frac{12}{\sqrt{MN}}\int_{\delta}^{R}\sqrt{\log 2\mathcal{N}(\mathcal{F},\epsilon,\|\cdot\|_{N,M,2})}\mathrm{d}\epsilon\right\}.$
@@
$\hat{\mathcal{R}}_{N,M}(\mathcal{F})$ $\leq 4R\sqrt{\frac{d}{MN}}+12R\sqrt{\frac{1}{MN}\log 2\mathcal{N}(\mathcal{F},R\sqrt{d(MN)^{-1}},\|\cdot\|_{N,M,\infty})}$
@@
For $\mathbf{x}_{1},\ldots,\mathbf{x}_{M}\in\mathbb{R}^{dN}$, we set $g(\mathbf{x}_{1},\ldots,\mathbf{x}_{M})=\sup_{z\in\mathcal{Z}}\left|\frac{1}{M}\sum_{j=1}^{M}\mathbb{E}_{X\sim\rho_{\mathbf{x}_{j}}}[h(X,z)]-\mathbb{E}_{X\sim\mu_{*}}[h(X,z)]\right|$. By the variational formulation of KL-divergence (e.g., Corollary 4.15 in Boucheron et al., (2013)), we get
@@
[…]
0: Dictionary sup_norm_dic maps NN to the average sup_norm
@@
  Initialize sup_norm_lst $\leftarrow[]$
@@
    $\text{{sup\_norm\_total}}\leftarrow\text{{sup\_norm\_total}}+\text{{sup\_norm}}$
@@
Figure 2: Averaged test ln(MSE) of singular MFNNs, across different NN and $\lambda$ for 5 epochs
Table 2: MSE comparison between merging $M=20$ networks across different NN and $\lambda$ after 5 epochs.
@@
Llama2 7B LoRA (best) 0 $80.55$ $82.86$ $83.19$ $81.60$ $71.08$ $84.51$ $71.90$ $90.21$ $80.74$
LoRA (ave.) 0 $64.73$ $76.31$ $77.76$ $68.70$ $57.02$ $69.02$ $69.04$ $70.63$ $69.15$
PoC merge 0 $32.29$ $62.57$ $83.58$ $22.20$ $28.41$ $29.42$ $61.53$ $28.50$ $43.56$
@@
Llama3 8B LoRA (best) 0 $80.45$ $88.47$ $86.82$ $87.60$ $82.25$ $90.87$ $73.85$ $95.78$ $85.76$
LoRA (ave.) 0 $80.51$ $88.87$ $86.85$ $87.00$ $80.78$ $90.98$ $73.71$ $95.84$ $85.57$
PoC merge 0 $81.73$ $88.96$ $87.77$ $88.00$ $81.40$ $91.71$ $74.46$ $96.45$ $86.31$
@@
Two layer mean-field neural networks provide a tractable analytical framework for studying infinitely wide neural networks. As $N\rightarrow\infty$, the optimization dynamics is captured by a partial differential equation (PDE) of the parameter distribution, where convexity can be exploited to show convergence to the global optimal solution (Chizat and Bach,, 2018; Mei et al.,, 2018; Rotskoff et al.,, 2019). If Gaussian noise is added to the gradient, we get MFLD which achieves global convergence to the optimal solution (Mei et al.,, 2018; Hu et al.,, 2019). Under the uniform LSI, Nitanda et al., (2022); Chizat, (2022) show that MFLD converges at an exponential rate by using the proximal Gibbs distribution associated with the dynamics. MFLD has attracted significant attention because of its feature learning capabilities (Suzuki et al., 2023b, ; Mousavi-Hosseini et al.,, 2024).
As the assumption that $N=\infty$ is not applicable to real-world scenarios, a discrete-time finite particle system would align closer to an implementable MFLD i.e. noisy gradient descent. Nitanda et al., (2022) provides a global convergence rate analysis for the discrete-time update by extending the one-step interpolation argument for Langevin dynamics (Vempala and Wibisono,, 2019). Meanwhile, the approximation error of the finite particle setting is studied in propagation of chaos literature (Sznitman,, 1991). For finite MFLD setting, Mei et al., (2018) first suggested that approximation error grows exponentially with time before Chen et al., (2022); Suzuki et al., 2023a proved the uniform-in-time propagation of chaos with error bound: $O\left(\frac{\alpha}{\lambda N}\right)$, suggesting that the difference between the finite $N$-particle system and mean-field limit shrinks as $N\rightarrow\infty$. However, this also means that particle approximation error blows-up exponentially as $\lambda\rightarrow 0$ due to the exponential relationship between $\alpha$ and $\lambda$ (Nitanda et al.,, 2022; Chizat,, 2022; Suzuki et al., 2023a, ). Suzuki et al., 2023b proposes an annealing procedure for classification tasks to remove this exponential dependence in LSI, wihch requires that $\lambda$ be gradually reduced over time and will not work for fixed regularization parameters. Nitanda, (2024); Chewi et al., (2024) then proved a refined propagation of chaos independent of $\alpha$ at the solution as described in Section 1.1.